# Structural Anisotropy Stabilises Asymmetric Beating in Instability Driven Filaments

### v1

@@ 4 Aniostropic filaments with no intrinsic curvature
Figure 6: (a) Results of numerical simulations of an anisotropic filament with no intrinsic curvature $(\kappa_{0}=0,\beta\geq 1)$. (b) Snapshots of filaments over one period undergoing whirling dynamics in the isotropic case, close to the bifurcation $(\beta=1,f=36)$, an anisotropic case near the bifurcation $(\beta=1.2,f=46)$ and an anisotropic case further from the bifurcation $(\beta=1.2,f=100)$.
Figure 7: (a) Bifurcation diagrams for an anisotropic filament with no intrinsic curvature ($\kappa_{0}=0$) for $\beta=1.2,1.3$ and $1.4$ (top to bottom). Dotted/full lines refer to unstable/stable solutions. (b) The real part of the four largest eigenvalues associated to the steady state for $\beta=1.5$. The eigenvectors allow us to identify the branches corresponding to in-plane beating (bottom left), and off-plane beating (bottom right). (c) The real part of the second largest eigenvalue associated to off-plane beating for various values of $\beta.$ We associate this eigenvalue becoming stable with P4 collapsing onto the off-plane beating branch. (d) The real part of the largest eigenvalue associated to the in-plane beating state for various values of $\beta$. For $\beta>\beta^{*}$, we see the eigenvalue remains stable.
@@ 4.1 Initial Value Problems
Now, we focus on the effect of introducing anisotropy through the bending modulus by introducing the parameter $\beta$, which describes the ratio of the bending moduli in the $\hat{\bm{\nu}}-$ and $\hat{\bm{\mu}}-$directions. Throughout this paper we focus on $\beta>1,$ but the $\beta<1$ case follows immediately by considering a rotation of the axes, and redefining the non-dimensional follower force as $\tilde{f}=\beta f$ (c.f. Eq (5)). In Figure 6a, we show the results of IVPs for $\beta\in[1,1.5]$ and $f\in[0,300].$ Regardless of $\beta$ and $f$, the steady state is the vertically upright filament. For $\beta=1$, this buckles into whirling, while for $\beta>1$ the steady state buckles into planar beating in the $(x,z)-$plane. For moderate values of $\beta$, beating becomes unstable as we increase the forcing, with whirling appearing in its place. Close to the bifurcation whirling is very similar to in-plane beating, with a small off-plane component. This causes the tip to trace an ellipse with a major axis that aligns with the $\hat{\bm{x}}-$axis, as shown in Figure 6b, middle and Supplemental Video 3. Further from the bifurcation, however, the off-plane component grows and whirling resembles the isotropic case more closely (see Figure 6b, right). For larger forcing values, whirling becomes unstable and planar beating in the $(x,z)-$plane reappears. As we increase $\beta$, the region in which whirling is stable appears to decrease, until $\beta=1.35$, where we no longer observe whirling through IVPs. For these $\beta$ values, we only observe planar beating in the $(x,z)-$plane after buckling.
@@ 4.2 Bifurcation and Stability Analyses
5 Using bending modulus anisotropy to stabilise asymmetric beating
@@
Figure 8: (a) Results of numerical simulations of an anisotropic filament with intrinsic curvature $(\kappa_{0}=3\pi/4,\beta\geq 1)$. (b) The eigenvalue associated to off-plane beating for several values of $\beta$. For $\beta\lesssim 1.4$, this eigenmode becomes unstable before the in-plane eigenmode (i.e. before $f=f^{*}$). For $\beta\gtrsim 1.4$, the in-plane eigenmode becomes unstable first. (c) The bifurcation diagram for $\kappa_{0}=3\pi/4$, $\beta=2.$ We see that only asymmetric planar beating (bottom) is stable.
@@
Following the previous section, we see that increasing $\beta>1$ has the effect of stabilising planar solutions. Thus we now combine the effects of $\kappa_{0}\neq 0$ and $\beta\neq 1$, with the goal of recovering the planar asymmetric beating observed in 2D simulations of filaments with intrinsic curvature, which were found to be unstable to 3D perturbations in Section 3 for $\beta=1$. We focus on $\kappa_{0}=3\pi/4$ to match previous studies [31], and increase $\beta$ from $\beta=1$ to $\beta=1.5.$ The results from IVPs over this parameter range are shown in Figure 8a.
@@
As discussed in previous sections, $\kappa_{0}$ and $\beta$ have opposite effects on the initial buckling modes; increasing the intrinsic curvature (i.e. increasing $\kappa_{0}$) causes the off-plane modes to buckle earlier, whereas increasing the anisotropy through the bending moduli (i.e. increasing $\beta$) causes later buckling of these modes. In both cases, in-plane buckling happens at the same critical forcing, regardless of $\kappa_{0}$ or $\beta$. Therefore, fixing $\kappa_{0}$ and increasing $\beta$ shifts the off-plane beating modes to buckle at higher values of the forcing. This is shown in Figure 8b, which displays the eigenvalues corresponding to off-plane buckling for various values of $\beta$. As we can see, at $\beta\approx 1.4$ the off-plane buckling occurs at $f=f^{*}$, the critical forcing at which in-plane buckling occurs. Beyond this $\beta,$ the off-plane buckling occurs for $f>f^{*}$, indicating that the initial buckling of the filament will be to in-plane beating, and will occur at $f=f^{*}.$ Accordingly, as we increase $\beta$ from $1$, the IVPs in Figure 8a show that the initial buckling into P1 happens at a critical value of the forcing which increases with $\beta$ until it reaches $f=f^{*}$ at $\beta\approx 1.4,$ beyond which the filament buckles into in-plane asymmetric beating at $f=f^{*}$.
@@
Furthermore, as increasing $\beta$ increases the critical forcing value at which off-plane beating bifurcates from the steady state, increasing the bending modulus anisotropy results in the birth of whirling again (as illustrated by the IVPs in Figure 8a). For these intermediate values of $\beta,$ Floquet analysis indicates bistability between P1, whirling and in-plane beating at several points, likely linked to the existence of distinct unstable quasiperiodic and periodic solutions joining the three branches. However as $\beta$ is increased further, and the off-plane buckling modes become unstable for larger $f,$ we reach a scenario where the whirling branch dies again above a critical value of $\beta$. This occurs through the same mechanism as the $\kappa_{0}=0$ and $\beta=1$ cases.
@@
As expected, increasing the bending modulus anisotropy further leads to a complete collapse onto planar dynamics. While initial value problems from the upright-state indicate that this collapse has occurred for $\beta=1.5$ (see Figure 8a), Floquet analysis shows that there are still regions where P1 is stable for this choice of parameters, alongside in-plane beating. By $\beta=2$, however only in-plane beating is stable, showcasing the same asymmetric beating dynamics as the planar case for $\kappa_{0}=3\pi/4$, with particular asymmetry close to the bifurcation. We show the bifurcation diagram for $\beta=2,$ and the planar beating dynamics, in Figure 8c and in Supplemental Video 4.
@@ 6 Summary and Discussion
In this paper, we investigated the effects of asymmetry and structural anisotropy in a follower-force driven filament, through the introduction of both intrinsic curvature and bending modulus anisotropy. In both cases, we observed that the initial buckling changed from a double Hopf bifurcation, which is observed in the symmetric case [37, 38], to a single Hopf, suggesting that the double Hopf is a special case of these systems and is related to symmetry. When varying only the intrinsic curvature, in 2D simulations we obtain the same asymmmetric beating reported in the literature [31]. However in 3D, we see that this beating is unstable, with the filament instead buckling into a fully 3D beating occurring primarily perpendicular to the plane of intrinsic curvature (P1). Increasing the forcing, we observe a tilted whirling and discuss how this changes with both $f$ and $\kappa_{0}$. For moderate curvatures, we observe that P1 and whirling become bistable, making the stable QP1 branch connecting whirling to beating in the zero-curvature case now unstable, and introducing further unstable solutions which we identify through bisection. Similar results hold when varying the bending modulus anisotropy, replacing P1 with planar beating in the plane of preferential bending. In both cases, we observe that as we increase the amount of asymmetry, whirling does not only become unstable, but rather vanishes completely as a solution. Instead, we observe only P1 for the intrinsic curvature case, or in-plane beating for the anisotropic bending modulus case. Finally, we show how introducing anisotropy through the bending modulus on a filament with intrinsic curvature can re-introduce whirling as a solution for moderate $\beta$ values, and ultimately stabilise the planar asymmetric beating observed in 2D for isotropic filaments with intrinsic curvature for larger $\beta$ values. We find that a difference by a factor of 2 in the bending moduli is sufficient for this collapse into planar beating for the highest value of $\kappa_{0}$ tested for the follower force model; a remarkably reasonable ratio.
@@
Changing the intrinsic curvature has the effect of adding asymmetry into the beating pattern, motivating its inclusion in recent numerical studies [26, 27, 28, 29, 30, 31]. These 2D works use values $\kappa_{0}\in[2,2.71]$ to match the beating pattern to experiments, in particular to wild-Chlamydomonas flagella [26, 29]. However this is not the only potential cause of asymmetry in ciliary beating; some works suggest that the bending modulus in each plane may not be constant, instead utilising a curvature dependent bending modulus [53, 23]. In such a model, the in-plane bending modulus can take one of two values depending on the sign of the local curvature. Other works suggest that the asymmetry is a result of bias in the rates at which the dynein attach and detach from the microtubules, depending on their placement in the axoneme [26, 27, 28]. In each of these models, numerical studies result in asymmetric dynamics exhibiting distinct recovery and effective strokes. Our results indicate that these asymmetric beating solutions could be unstable in 3D, but that they could be stabilised by incorporating anisotropy to the model, for instance structural anisotropy through the bending modulus.
@@
Incorporating anisotropy through the bending modulus in either direction is motivated by the internal structure of the cilium. Cilia has a preferential bending direction both due to the orientation of the central pair of microtubules [5, 20, 21] and the stiffness of the 5-6 bridge [5, 22] which limits sliding between microtubule doublets 5 and 6 in the axoneme. The effect of the bridge is captured in our work by allowing the perpendicular stiffness to differ from the in-plane stiffness, where ‘in-plane’ is aligned with the preferential beating plane of the cilium. Structural anisotropy was utilised in [24], where it was shown that the transition from planar to whirling dynamics is delayed as we increase the anisotropy. Based on our results, we suspect that increasing the anisotropy further in this study would eradicate the 3D whirling dynamics completely, and only planar solutions would remain. It is expected that larger values of $\beta$ may be required for different forcing mechanisms. For example, in [24], values up to $\beta=10$ were considered and 3D dynamics were still observed, while we found values as small as $\beta=2$ were sufficient to suppress 3D dynamics entirely for the follower force model. In [54], experimental evidence on sand dollar sperm flagella reports a value of $\beta=13.8.$ This is much higher than the threshold for entirely planar dynamics that we find, and so could explain the planar beating observed by these cells [54]. In [49], a wooden axonemal model supplemented with a stronger 5-6 bridge (by using a silicon adhesive between doublets 5 and 6) and a central pair was used to generate an estimate of $\beta=2.6$ for the axoneme.
@@
Although the follower force model differs from the internal shear-driven sliding mechanism which generates ciliary beating, the dynamics observed in this paper are reminiscient of those observed biologically. For instance nodal cilia found in the embryo are known to whirl with a tilt to the left, generating fluid flows which break the left-right symmetry in the embryo [16, 17, 18], similar to the asymmetric whirling observed for moderate $\kappa_{0}$ and $\beta$. Motile cilia on the other hand, such as respiratory cilia, undergo asymmetric near-planar beating with low-curvature power strokes, followed by distinct high-curvature recovery strokes [14]. While we do not observe these two components of the beat with this simple model, for large enough $\beta$ and moderate $\kappa_{0}$, we do observe asymmetric planar beating. Structurally, these cilia are different; nodal cilia lack the central pair of MTs in the axoneme, which respiratory cilia do have. The existence of a central pair would increase the structural anisotropy, and hence increase $\beta$. This is consistent with our findings that filaments with lower $\beta$ values can access tilted whirling as a state, whereas higher $\beta$-filaments can only achieve planar beating.
@@
The follower force model is a simple, commonly-used model that yields cilia-like dynamics through instabilities and provides an environment to explore the effect of anisotropies on these states. Some other forcing mechanisms more accurately describe the shear-driven activation of cilia, instead considering the forcing with a multi-level approach; considering how microscale attachment and detachment of dynein motors in the axoneme can lead to large scale ciliary beating. In such models, anisotropy can be included through the bending modulus [24], or by including stiff springs to represent the 5-6 bridge in the axoneme [23]. Then asymmetric beating can be achieved through the intrinsic curvature [29, 30], or through biased feedback mechanisms which control the forcing output of the motors [26, 27, 28]. Exploring how anisotropies through the curvature and stiffness affect dynamics in models with more accurate forcing mechanisms would give more accurate estimates for the $\kappa_{0}$ and $\beta$ values necessary to recover beats closer to those observed in the literature, and also perhaps give more knowledge of the core ingredients needed to obtain such beats. Another future avenue of interest is considering the collective motion of asymmetrically beating filaments, investigating if their hydrodynamic interactions can lead to synchronisation and perhaps even metachronal waves as observed experimentally. This is the subject of our future research.

### v2

@@ 4.1 Initial Value Problems
In the previous section we investigated the effect of intrinsic curvature defined in the plane defined by $\hat{\bm{\nu}}$ and $\hat{\bm{t}}$. Now, we set $\kappa_{0}=0$ and focus on the effect of introducing anisotropy through the bending stiffness by introducing the parameter $\beta$, which describes the ratio of the bending stiffnesses about the $\hat{\bm{\nu}}-$ and $\hat{\bm{\mu}}-$axes. Throughout this paper we focus on $\beta>1,$ but the $\beta<1$ case follows immediately by considering a rotation of the axes, and redefining the non-dimensional follower force as $\tilde{f}=\beta f$ (c.f. Eq (5)). In Figure 6a, we show the results of IVPs for $\beta\in[1,1.5]$ and $f\in[0,300].$ Regardless of $\beta$ and $f$, the steady state is the vertically upright filament. For $\beta=1$, this buckles into whirling, while for $\beta>1$ the steady state buckles into planar beating in the $(x,z)$-plane. For moderate values of $\beta$, beating becomes unstable as we increase the forcing, with whirling appearing in its place. The frequency of these whirling solutions for different $\beta$ and $f$ values can be found in Appendix C. Close to the bifurcation in which planar beating appears to become unstable, whirling is very similar to in-plane beating, with a small off-plane component. This causes the tip to trace an ellipse with a major axis that aligns with the $\hat{\bm{x}}-$axis, as shown in Figure 6b, middle and Supplemental Video 3. The large amplitude in the $\hat{\bm{x}}-$direction close to the bifurcation suggests that we can expect the whirling solution to branch from the planar beating solution branch, instead of the steady state. Further from the bifurcation, however, the off-plane component grows and whirling resembles the isotropic case more closely (see Figure 6b, right). For larger forcing values, whirling becomes unstable and planar beating in the $(x,z)$-plane reappears. As we increase $\beta$, the region in which whirling is stable appears to decrease, until $\beta=1.35$, where we no longer observe whirling through IVPs. For these $\beta$ values, we only observe planar beating in the $(x,z)$-plane after buckling.
Figure 6: (a) Results of numerical simulations of an anisotropic filament with no intrinsic curvature $(\kappa_{0}=0,\beta\geq 1)$. The black lines are drawn to indicate the approximate boundaries of the solution regions. (b) Snapshots of filaments over one period undergoing whirling dynamics in the isotropic case, close to the bifurcation $(\beta=1,f=36)$, an anisotropic case near the bifurcation $(\beta=1.2,f=46)$ and an anisotropic case further from the bifurcation $(\beta=1.2,f=100)$. The three dynamics are indicated on the phase diagram in (a) using numbered circles.
Figure 7: (a) Bifurcation diagrams showing the different solutions as we vary $f$, and the corresponding mean bending energy of these solutions, $\bar{E}_{b}$, for an anisotropic filament with no intrinsic curvature ($\kappa_{0}=0$) for $\beta=1.2,1.3$ and $1.4$ (top to bottom). Dashed/full lines refer to unstable/stable solutions. Square markers are used to indicate the location of bifurcations as we vary $f$. We see the separation of the in-plane and off-plane solution branches as $\beta$ increases, and observe bistability between in-plane beating and whirling (inset). For larger values of $\beta$, the whirling solution branch vanishes completely (bottom). (b) The real part of the four largest eigenvalues associated to the steady state for $\beta=1.5$. As described in the main text, we visualise the eigenvectors, $\bm{x}_{i}$ for i=1,,4i=1,...,4, by taking linear combinations of the complex conjugate pairs, for example $\bm{v}_{1}=a(\bm{x}_{1}+\bm{x}_{2})$ where $a$ is a constant. We plot the eigenvectors for $a=5$ to identify the branches corresponding to in-plane beating (bottom left), and off-plane beating (bottom right). (c) The real part of the second largest eigenvalue associated to off-plane beating for various values of $\beta.$ We associate this eigenvalue becoming stable with P4 collapsing onto the off-plane beating branch. (d) The real part of the largest eigenvalue associated to the in-plane beating state for various values of $\beta$. For $\beta>\beta^{*}$, we see the eigenvalue remains stable.
@@ 4.2 Bifurcation and Stability Analyses
5 Using bending stiffness anisotropy to stabilise asymmetric beating
@@
Figure 8: (a) Results of numerical simulations of an anisotropic filament with intrinsic curvature $(\kappa_{0}=3\pi/4,\beta\geq 1)$. The black lines are drawn to indicate the approximate boundaries of the solution regions. (b) The eigenvalue associated to off-plane beating for several values of $\beta$. For $\beta\lesssim 1.4$, this eigenmode becomes unstable before the in-plane eigenmode (i.e. before $f=f^{*}$). For $\beta\gtrsim 1.4$, the in-plane eigenmode becomes unstable first. (c) The bifurcation diagram showing the different solutions as we vary $f$, and the corresponding mean bending energy of these solutions, $\bar{E}_{b}$, for $\kappa_{0}=3\pi/4$, $\beta=2.$ We see that only asymmetric planar beating (bottom) is stable.
@@
Following the previous section, we see that increasing $\beta>1$ has the effect of stabilising planar solutions. Thus we now combine the effects of $\kappa_{0}\neq 0$ and $\beta\neq 1$, with the goal of recovering the planar asymmetric beating observed in 2D simulations of filaments with intrinsic curvature, which were found to be unstable to 3D perturbations in Section 3 for $\beta=1$. Both the P1 and whirling states must be eradicated for a collapse onto planar dynamics. Following the results of the previous sections, we anticipate that the stability of the P1 branch and existence of whirling solution are related to the eigenmodes of the initial buckling event. As increasing $\kappa_{0}$ appears to cause an earlier buckling of the off-plane eigenmodes, while increasing $\beta$ delays the onset of this instability, we expect that increasing the intrinsic curvature will require a larger amount of anisotropy to stabilise the planar modes. As such, we focus on $\kappa_{0}=3\pi/4$, anticipating this to provide an upper bound for the value of $\beta$ which causes a collapse to 2D dynamics, and increase $\beta$ from $\beta=1$ to $\beta=1.5.$ The results from IVPs over this parameter range are shown in Figure 8a.
@@
As discussed in previous sections, $\kappa_{0}$ and $\beta$ have opposite effects on the initial buckling modes; increasing the intrinsic curvature (i.e. increasing $\kappa_{0}$) causes the off-plane modes to buckle earlier, whereas increasing the anisotropy through the bending stiffness ratio (i.e. increasing $\beta$) causes later buckling of these modes. In both cases, in-plane buckling happens at the same critical forcing, regardless of $\kappa_{0}$ or $\beta$. Therefore, fixing $\kappa_{0}$ and increasing $\beta$ shifts the off-plane beating modes to buckle at higher values of the forcing. This is shown in Figure 8b, which displays the eigenvalues corresponding to off-plane buckling for various values of $\beta$. As we can see, at $\beta\approx 1.4$ the off-plane buckling occurs at $f\approx f^{*}$, the critical forcing at which in-plane buckling occurs. Beyond this $\beta,$ the off-plane buckling occurs for $f>f^{*}$, indicating that the initial buckling of the filament will be to in-plane beating, and will occur at $f\approx f^{*}.$ Accordingly, as we increase $\beta$ from $1$, the IVPs in Figure 8a show that the initial buckling into P1 happens at a critical value of the forcing which increases with $\beta$ until it reaches $f\approx f^{*}$ at $\beta\approx 1.4,$ beyond which the filament buckles into in-plane asymmetric beating at $f\approx f^{*}$.
@@
Furthermore, as increasing $\beta$ increases the critical forcing value at which off-plane beating bifurcates from the steady state, increasing the bending stiffness anisotropy results in the birth of whirling again (as illustrated by the IVPs in Figure 8a). For these intermediate values of $\beta,$ Floquet analysis indicates bistability between P1, whirling and in-plane beating at several points, likely linked to the existence of distinct unstable quasiperiodic and periodic solutions joining the three branches. However as $\beta$ is increased further, and the off-plane buckling modes become unstable for larger $f,$ we reach a scenario where the whirling branch dies again above a critical value of $\beta$. This occurs through the same mechanism as the $\kappa_{0}=0$ and $\beta=1$ cases.
@@
As expected, increasing the bending stiffness anisotropy further leads to a complete collapse onto planar dynamics. While initial value problems from the upright-state indicate that this collapse has occurred for $\beta=1.5$ (see Figure 8a), Floquet analysis shows that there are still regions where P1 is stable for this choice of parameters, alongside in-plane beating. By $\beta=2$, however only in-plane beating is stable, showcasing the same asymmetric beating dynamics as the planar case for $\kappa_{0}=3\pi/4$, with particular asymmetry close to the bifurcation. We show the bifurcation diagram for $\beta=2,$ and the planar beating dynamics, in Figure 8c and in Supplemental Video 4.
@@ 6 Summary and Discussion
In this paper, we investigated the effects of asymmetry and structural anisotropy in a follower-force driven filament, through the introduction of both intrinsic curvature and bending stiffness anisotropy. In both cases, we observed that the initial buckling changed from a double Hopf bifurcation, which is observed in the symmetric case (Clarke2024BifurcationsFilaments, ; Schnitzer2025OnsetFilament, ), to a single Hopf, suggesting that the double Hopf is a special case of these systems and is related to symmetry. When varying only the intrinsic curvature, in 2D simulations we obtain the same asymmmetric beating reported in the literature (Wang2023Generalized-NewtonianFilament, ). However in 3D, we see that this beating is unstable, with the filament instead buckling into a fully 3D beating occurring primarily perpendicular to the plane the filament would deform in the absence of the follower force (P1). Increasing the forcing, we observe a tilted whirling and discuss how this changes with both $f$ and $\kappa_{0}$. For moderate curvatures, we observe that P1 and whirling become bistable, making the stable QP1 branch connecting whirling to beating in the zero-curvature case now unstable, and introducing further unstable solutions which we identify through bisection. Similar results hold when varying the bending stiffness anisotropy, replacing P1 with planar beating in the plane of preferential bending. In both cases, we observe that as we increase the amount of asymmetry, whirling does not only become unstable, but rather vanishes completely as a solution. Instead, we observe only P1 for the intrinsic curvature case, or in-plane beating for the anisotropic bending stiffness case. Finally, we show how introducing anisotropy through the bending stiffness on a filament with intrinsic curvature can re-introduce whirling as a solution for moderate $\beta$ values, and ultimately stabilise the planar asymmetric beating observed in 2D for isotropic filaments with intrinsic curvature for larger $\beta$ values. We find that a difference by a factor of 2 in the bending stiffnesses is sufficient for this collapse into planar beating for the highest value of $\kappa_{0}$ tested for the follower force model; a remarkably reasonable ratio.
@@
Changing the intrinsic curvature has the effect of adding asymmetry into the beating pattern, motivating its inclusion in recent numerical studies (Chakrabarti2019SpontaneousMicrofilaments, ; Chakrabarti2019HydrodynamicFilaments, ; Chakrabarti2022ACilia, ; Sartori2016DynamicFlagella, ; Cass2023TheFlagella, ; Wang2023Generalized-NewtonianFilament, ). These 2D works use values $\kappa_{0}\in[2,2.71]$ to match the beating pattern to experiments, in particular to wild-Chlamydomonas flagella (Chakrabarti2019SpontaneousMicrofilaments, ; Sartori2016DynamicFlagella, ). However this is not the only potential cause of asymmetry in ciliary beating; some works suggest that the bending stiffness in each plane may not be constant, instead utilising a curvature dependent bending stiffness (Moreau2024MinimalCilium, ; Han2018SpontaneousCilia, ). In such a model, the in-plane bending stiffness can take one of two values depending on the sign of the local curvature. Other works suggest that the asymmetry is a result of bias in the rates at which the dynein attach and detach from the microtubules, depending on their placement in the axoneme (Chakrabarti2019SpontaneousMicrofilaments, ; Chakrabarti2019HydrodynamicFilaments, ; Chakrabarti2022ACilia, ). In each of these models, numerical studies result in asymmetric dynamics exhibiting distinct recovery and effective strokes. Our results indicate that these asymmetric beating solutions could be unstable in 3D, but that they could be stabilised by incorporating anisotropy to the model, for instance structural anisotropy through the bending stiffness.
@@
Incorporating anisotropy through the bending stiffness in either direction is motivated by the internal structure of the cilium. Cilia have a preferential bending direction both due to the orientation of the central pair of microtubules (Gibbons1981CiliaEukaryotes, ; Marshall2006Cilia:Antenna, ; Gilpin2020TheFlagella, ) and the stiffness of the 5-6 bridge (Gibbons1981CiliaEukaryotes, ; AFZELIUS1959ElectronFixative., ) which limits sliding between microtubule doublets 5 and 6 in the axoneme. The effect of the bridge is captured in our work by allowing the perpendicular stiffness to differ from the in-plane stiffness, where ‘in-plane’ is aligned with the preferential beating plane of the cilium. Structural anisotropy was utilised in (Rallabandi2022Self-sustainedFlagellum, ), where it was shown that the transition from planar to whirling dynamics is delayed as we increase the anisotropy. Based on our results, we suspect that increasing the anisotropy further in this study would eradicate the 3D whirling dynamics completely, and only planar solutions would remain. It is expected that larger values of $\beta$ may be required for different forcing mechanisms. For example, in (Rallabandi2022Self-sustainedFlagellum, ), values up to $\beta=10$ were considered and 3D dynamics were still observed, while we found values as small as $\beta=2$ were sufficient to suppress 3D dynamics entirely for the follower force model. In (Ishijima1994FlexuralFlagella, ), experimental evidence on sand dollar sperm flagella reports a value of $\beta=13.8.$ This is much higher than the threshold for entirely planar dynamics that we find, and so could explain the planar beating observed by these cells (Ishijima1994FlexuralFlagella, ). In (Lindemann2016FunctionalFlagellum, ), a wooden axonemal model supplemented with a stronger 5-6 bridge (by using a silicon adhesive between doublets 5 and 6) and a central pair was used to generate an estimate of $\beta=2.6$ for the axoneme.
@@
Although the follower force model differs from the internal shear-driven sliding mechanism which generates ciliary beating, the dynamics observed in this paper are reminiscent of those observed biologically. For instance nodal cilia found in the embryo are known to whirl with a tilt to the left, generating fluid flows which break the left-right symmetry in the embryo (Nonaka1998RandomizationProtein, ; Smith2008FluidCilia, ; Smith2011MathematicalCilia, ), similar to the asymmetric whirling observed for moderate $\kappa_{0}$ and $\beta$. Motile cilia on the other hand, such as respiratory cilia, undergo asymmetric near-planar beating with low-curvature power strokes, followed by distinct high-curvature recovery strokes (Chilvers2000AnalysisMethods, ). While we do not observe these two components of the beat with this simple model, for large enough $\beta$ and moderate $\kappa_{0}$, we do observe asymmetric planar beating. Structurally, these cilia are different; nodal cilia lack the central pair of MTs in the axoneme, which respiratory cilia do have. The existence of a central pair would increase the structural anisotropy, and hence increase $\beta$. This is consistent with our findings that filaments with lower $\beta$ values can access tilted whirling as a state, whereas higher $\beta$-filaments can only achieve planar beating.
@@
The follower force model is a simple, commonly-used model that yields cilia-like dynamics through instabilities and provides an environment to explore the effect of anisotropies on these states. Some other forcing mechanisms more accurately describe the shear-driven activation of cilia, instead considering the forcing with a multi-level approach; considering how microscale attachment and detachment of dynein motors in the axoneme can lead to large scale ciliary beating. In such models, anisotropy can be included through the bending stiffness (Rallabandi2022Self-sustainedFlagellum, ), or by including stiff springs to represent the 5-6 bridge in the axoneme (Han2018SpontaneousCilia, ). Then asymmetric beating can be achieved through the intrinsic curvature (Sartori2016DynamicFlagella, ; Cass2023TheFlagella, ), or through biased feedback mechanisms which control the forcing output of the motors (Chakrabarti2019SpontaneousMicrofilaments, ; Chakrabarti2019HydrodynamicFilaments, ; Chakrabarti2022ACilia, ). Exploring how anisotropies through the curvature and stiffness affect dynamics in models with more accurate forcing mechanisms would give more accurate estimates for the $\kappa_{0}$ and $\beta$ values necessary to recover beats closer to those observed in the literature, and also perhaps give more knowledge of the core ingredients needed to obtain such beats.
[…]
stiffnessThe anisotropies studied in this work are assumed to be constant along the filament length, but in reality this may not be the case. Towards the bottom of each cilium, connecting the axoneme to the basal body, is a region called the transition zone, in which several structural changes from the ‘9+2’ axoneme can be found (Mercey2024TheZones, ). Incorporating the structural properties of this region is of interest, as defective proteins in the transition zone are known to cause a range of human diseases, called ciliopathies (Goncalves2017TheGate, ). Investigating the role of this coupling at the base, either through different boundary conditions at the wall or through spatially-varying anisotropies, will provide an insight to the impact of these localised anisotropies on ciliary dynamics. Another future avenue of interest is considering the collective motion of asymmetrically beating filaments, stiffnesssuch as in ctenophores, which use plates of synchronously beating cilia for locomotion (Tamm2014CiliaCtenophores, ). The plates of cilia coordinate such that a constant phase difference is observed between neighbouring plates, causing a metachronal wave along the ctenophore length. Each cilium has a ‘9+3’ axoneme (Afzelius1961TheSwimming-plates., ), suggesting anisotropy at the individual level is important to ctenophore motility. However it is not immediately clear how the structure of these comb plates, in particular how each cilium is bound to the plate, will affect both the individual and collective dynamics. The study of collective dynamics, in particular investigating if hydrodynamic interactions between cilia can lead to synchronisation and perhaps even metachronal waves as observed experimentally, will be the subject of our future research.
Declaration of Interests
The authors report no conflict of interest.
Acknowledgement
B.C. gratefully acknowledges funding from an EPSRC scholarship (Grant No. EP/W523872/1).